OpenStudy (anonymous):

i need help on this question...let f(X)=8x+2, find the average rate of change from 5 to 8

OpenStudy (anonymous):

sorry the denominator was wrong. \[\frac{f(8)-f(5)}{8-5}\]

OpenStudy (anonymous):

\[f(8)=64+2=66\] \[f(5)=25+2=27\] \[8-5=3\] \[\frac{66-27}{3}=\frac{41}{3}\]

OpenStudy (anonymous):

so the answer is 43/3?

OpenStudy (anonymous):

no i messed up on the last answer. is should be \[\frac{66-27}{3}=\frac{39}{3}=13\]

OpenStudy (anonymous):

ok, thanks alot, can i ask you another question?